A final exam in statistic has a mean of 73 with a standard deviation of 7.73. Random sample of 24 students is selected and the mean test score of the sample is computed. What percentage of sample means are less than 70?
Answer by stanbon(75887)

A final exam in statistic has a mean of 73 with a standard deviation of 7.73. Random sample of 24 students is selected and the mean test score of the sample is computed. What percentage of sample means are less than 70?
-----------------------
z(70) = (70-73)/[7.73/sqrt(24)] = 1.9013
----
P(x-bar < 70) = P(z < 1.9013) = 0.9714
=================================
Cheers,
Stan H.
===============